# arXiv:2603.01692v1[cs.LG] 02 Mar 2026

### v1

@@ 3.4 Structured Reasoning
After validation and memory update, the reasoning module generates the next improvement hypothesis, serving as the gradient signal that directs optimization.
Challenge Identification and Hypothesis Generation.
The module extracts challenges from structured feedback using the trace’s local history:
$\mathcal{C}^{(i)}_{t}=\texttt{ExtractChallenges}(f^{(i)}_{t},\mathcal{H}^{(i)},\texttt{desc}),$
where $\mathcal{H}^{(i)}$ is the trace’s local experiment history and desc is the task description. The balance shifts from exploratory analysis (identifying gaps relative to SOTA practices) toward targeted diagnosis (addressing specific issues revealed by execution feedback). For each identified challenge $c\in\mathcal{C}^{(i)}_{t}$, the module generates a concrete hypothesis:
$\eta_{c}=\texttt{Hypothesize}(c,s^{*(i)},f^{(i)}_{t},\mathcal{H}^{(i)}),$ (7)
producing candidate directions $\{(c,\eta_{c})\}$ based on local information.
Prioritization and Selection.
[…]
@@
where the success memory $\mathcal{M}$3.3) modulates scoring: hypotheses resembling past successes gain confidence, while those similar to past failures receive penalties. Rather than greedily selecting the top-scored hypothesis, we sample from top-$k$ candidates to maintain exploration diversity, yielding the selected hypothesis $\eta^{(i)}_{t+1}$. In multi-trace settings, additional cross-trace mechanisms further guide selection (§3.5).
Solution Update.
The selected hypothesis specifies both the target component and modification strategy:
$s^{(i)}_{t+1}=\texttt{Implement}(\eta^{(i)}_{t+1},s^{*(i)}),$ (9)
where implementation generates a code sketch from the hypothesis, then iteratively refines until successful execution (§3.6).
@@ 3.5 Multi-trace Optimization
Single-trace optimization risks converging to local optima and cannot recover from poor early decisions. Multi-trace parallelism addresses these limitations by maintaining $N$ traces that evolve independently while sharing knowledge through success memory. The success memory $\mathcal{M}$3.3) enables online knowledge sharing: when trace $i$ discovers an effective strategy, that strategy becomes immediately available to traces $j\neq i$, unlike parallel search methods that maintain independent trajectories.
@@
At initialization, we enforce diversity through sequential hypothesis generation:
$\eta_{0}^{(n)}=\texttt{GenerateHypothesis}\bigl(\mathcal{T},\{\eta_{0}^{(j)}\}_{j<n}\bigr),$
Initial hypotheses are sequentially conditioned on prior proposals to ensure complementary exploration, and implemented as $s_{0}^{(n)}=\texttt{Implement}(\eta_{0}^{(n)},\mathcal{T})$. Subsequent execution divergence naturally maintains diversity.
@@
where $\{f_{\eta}\}$ denotes the associated feedback stored alongside each hypothesis. The selector has flexibility to select, modify, or generate based on the candidate pool and historical outcomes.
@@ 4 Experiments
We evaluate Gome on MLE-Bench under a rigorous closed-world protocol to validate our claim that gradient-based optimization scales more effectively than search as reasoning capability improves. Additional experiments including protocol comparison and cost analysis are provided in Appendix A.
@@ 4.1 Experiment Setup
All experiments are conducted on MLE-Bench (Chan et al., 2024), a comprehensive benchmark comprising 75 Kaggle competitions for evaluating AI agents on machine learning engineering tasks, categorized by complexity into low, medium, and high. MLE-Bench-Lite consists of 22 low-complexity tasks for efficient evaluation. We use the official metrics with any-medal rate as the primary indicator.
Implementation details.
We evaluate Gome using three frontier LLMs representing a spectrum of reasoning capabilities: DeepSeek-R1 (Guo et al., 2025a), o3 (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a), ordered by increasing reasoning strength. For the test environment, we use 12 vCPUs, 220GB RAM, and 1 NVIDIA V100 GPU with a 12-hour time budget. All results are averaged over three runs with different random seeds.
Methods for comparison.
We compare Gome against leading closed-world MLE agents on MLE-Bench, including MLAB (Huang et al., 2023), OpenHands (Wang et al., 2024), AIDE (Jiang et al., 2025), AIRA (Toledo et al., 2025), and ML-Master (Liu et al., 2025). For ML-Master, we re-evaluate under identical hardware and time budget; for other baselines, we use results reported in their respective papers or the official leaderboard.
@@ 4.2 Main Results
Gome establishes a new state-of-the-art on MLE-Bench. As shown in Table 3, Gome outperforms all closed-world baselines, achieving a peak any-medal rate of 35.1% with GPT-5. Beyond aggregate performance, Gome excels in solution quality: it achieves a 96.0% valid submission rate and converts 16.4% of tasks to Gold medals. A critical trend emerges when comparing against the previous gradient-free state-of-the-art, ML-Master, under identical constraints (12h, V100). While performance is comparable on weaker reasoners (DeepSeek-R1: 23.4% vs. 22.7%), Gome pulls ahead significantly with stronger models, widening the gap to +11.1% with GPT-5 and substantially improving Gold rate (16.4% vs. 12.4%). This widening gap provides initial evidence for our claim: gradient-based optimization scales more effectively with reasoning capability, a trend we analyze systematically in Section 5. Gome also demonstrates superior efficiency: even with the same o3 backbone, Gome surpasses AIRA’s medal rate (32.5% vs. 31.6%) while matching its solution quality (Median+ and Gold), despite operating with half the time budget (12h vs. 24h) and weaker hardware (V100 vs. H200).
@@
Accurate gradient signals drive rapid convergence. On MLE-Bench-Lite (Table 4), Gome achieves a 68.2% medal rate, matching the best open-world method (Leeroo) despite lacking external retrieval (see Appendix A.3 for open-world comparison). This validates the efficiency of the gradient-based paradigm: on low-to-medium complexity tasks, diagnostic feedback acts as a high-fidelity gradient, enabling rapid convergence analogous to descent on well-conditioned problems (Nesterov and others, 2018). Conversely, the performance plateau on high-complexity tasks across all methods echoes our core claim regarding signal quality: as problem difficulty increases, diagnostic reasoning becomes inherently harder (yielding “noisier gradients”), suggesting that further gains require either stronger reasoners or external knowledge to restore signal accuracy.
4.3 Ablation Study
To assess the contributions of Gome’s core components, we conduct an ablation study comparing the full model against three variants: (1) w/o Structured Reasoning, replacing diagnostic analysis with simple error-correction prompts; (2) w/o Success Memory, removing the global repository of proven patterns; and (3) w/o Multi-trace Optimization, reducing to single-trajectory optimization without cross-trace knowledge sharing. As shown in Table 5, removing any component degrades both final outcomes (Medal, Gold) and optimization quality (Impr., IC).
@@ 4.3 Ablation Study
The structured reasoning ablation causes the most severe degradation, with improvement rate dropping from 41.1% to 22.6%, indicating that without diagnostic analysis, most optimization steps fail to produce valid improvements. Removing success memory shows a different failure mode: while per-iteration success remains reasonable (36.2%), the lack of momentum from proven patterns leads to redundant exploration, reducing medal rate by 6.2%. In contrast, multi-trace ablation maintains comparable improvement rate but achieves lower final performance, suggesting that cross-trace knowledge sharing primarily helps escape local optima. These results confirm that all three components (Table 2) contribute meaningfully to Gome’s performance.
@@ 5 Scaling Analysis
Figure 3: Gome performance scales with model reasoning capability. As base model capability increases from Efficiency to Frontier tiers, Gome’s advantage over search-based baselines widens significantly.
Figure 4: Optimization dynamics across model capability tiers. Gome (solid red) exhibits rapid early convergence, while MCTS (dashed blue) starts slower but catches up on weaker models. On Frontier-tier models, Gome maintains its advantage throughout.
To systematically validate our claim that gradient signal quality depends on reasoning capability, we evaluate Gome across three tiers: Efficiency (e.g., GPT-4o-mini), Advanced (e.g., DeepSeek-R1), and Frontier (e.g., o3, GPT-5) [^3] See Appendix A.1 for complete tier definitions and per-model results.. We compare against Gome-MCTS, a controlled variant replacing gradient updates with tree search within identical prompt and infrastructure, isolating the impact of the optimization strategy.
Figure 3 reveals a distinct phase transition in final performance. On Efficiency-tier models, Gome lags behind search-based baselines: when gradient signals are noisy due to weak reasoning, exhaustive search outperforms directed updates. However, a crossover emerges as capability increases: search-based gains plateau while Gome accelerates, widening the gap to +5.8% on o3 and +7.1% on GPT-5. Figure 4 further reveals how this gap emerges over time. Gome exhibits rapid early convergence across all tiers, but the divergence lies in sustainability: on weaker models (GPT-4o), unreliable gradients cause premature convergence, allowing MCTS to eventually catch up; on Frontier models, accurate gradient signals enable Gome to maintain its lead throughout, achieving both faster convergence and superior final performance.
@@ 6 Conclusion
We presented Gome, a framework that reframes MLE agents through gradient-based optimization. By treating LLM reasoning as gradient signals, success memory as momentum, and multi-trace execution as distributed optimization, Gome establishes a principled correspondence between classical optimization and agentic ML engineering.
Our experiments demonstrate state-of-the-art performance on MLE-Bench, while scaling analysis reveals a more fundamental finding: tree search scales with inference compute through exhaustive exploration, while Gome scales with model capability through increasingly accurate gradient signals. As foundation models advance in reasoning ability, this positions gradient-based optimization as an increasingly favorable paradigm.
This work opens a new design axis for MLE agents: rather than engineering more sophisticated search strategies, future systems may benefit from investing in gradient quality through better diagnostic reasoning and richer feedback signals. We release our codebase and GPT-5 traces at https://anonymous.4open.science/r/GOME-44E0.
@@
Gome’s effectiveness is sensitive to the base model’s reasoning capability. As shown in our scaling analysis, Gome underperforms tree-search methods with efficiency-tier models, where weaker reasoners may hallucinate diagnostic feedback. This also implies higher inference costs when using frontier models. However, this dependency aligns with the trajectory of LLM development: as reasoning-oriented models advance rapidly, the capability threshold for effective gradient-based optimization will become increasingly accessible.
@@
We evaluated under a closed-world protocol on MLE-Bench, restricting agents to task-provided materials without external knowledge retrieval. While this ensures rigorous comparison of optimization strategies by controlling for confounding factors such as retrieval quality and knowledge source coverage, it limits generalizability: real-world MLE tasks often benefit from external knowledge, and Gome’s performance in open-world settings remains less thoroughly validated (see Appendix A.3 for preliminary open-world results). Additionally, real Kaggle competitions allow participants to access shared notebooks and public baselines, so our closed-world setting may underestimate performance achievable in practical scenarios.
@@
Our gradient-based framing establishes functional correspondence rather than mathematical equivalence. Unlike true gradients in continuous optimization, LLM-generated update signals lack formal properties such as convergence guarantees. This means that in rugged landscapes where solutions require fundamentally different paradigms rather than incremental edits, gradient-based guidance may get trapped in local optima. In practice, this is mitigated by the nature of MLE tasks: ML pipelines are typically repairable, making the smoothness assumption reasonable for most practical scenarios. Our multi-trace optimization and success memory further address this by enabling shared exploration across trajectories.
@@ A.1 Extended Scaling Analysis
This section provides detailed per-model results supporting the scaling analysis in Section 5. We categorize models into three capability tiers following the stratification established by the Artificial Analysis Intelligence Index [^4] https://artificialanalysis.ai/, SWE-bench, and Aider leaderboard [^5] https://aider.chat/docs/leaderboards/ (Artificial Analysis Team, 2025; Jimenez et al., 2023; Gauthier, 2025): Efficiency models are optimized for throughput and cost; Advanced models represent balanced capability-efficiency trade-offs; Frontier models define current state-of-the-art reasoning.
@@ A.5 Task-type Breakdown
Table 10 presents detailed score comparison across 9 representative MLE-Bench tasks, grouped by data modality. We evaluate Gome with three frontier-tier backbone models: DeepSeek-R1 (Guo et al., 2025a), o3 (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a).
@@ B.1 Gome Hyperparameters
B.2 ML-Master Reproduction
We reproduce ML-Master (Liu et al., 2025) as a representative state-of-the-art tree search baseline. We obtain the official implementation from the authors’ public repository [^6] https://github.com/sjtu-sai-agents/ML-Master and adapt it to our evaluation protocol.
@@ B.2 ML-Master Reproduction
Table 12: ML-Master reproduction validation on DeepSeek-R1. Original results are from Liu et al. (2025) on A100 GPUs; ours are averaged over 3 independent runs under closed-world protocol on V100 GPUs. All values are any-medal rates (%) on MLE-Bench with 12-hour budget.
@@ B.3 Gome-MCTS Implementation
Our modular framework naturally supports diverse exploration strategies beyond gradient-based optimization. To provide a fair comparison between optimization paradigms within the same framework, we implement Monte Carlo Tree Search (MCTS), a widely-adopted algorithm in existing MLE agents such as ML-Master (Liu et al., 2025). This implementation, called Gome-MCTS, serves two purposes: (1) it validates that our framework is a general-purpose platform capable of accommodating different search algorithms, and (2) it isolates the effect of the optimization strategy from other architectural differences when comparing against external baselines.

### v2

@@ 3.4 Structured Reasoning
The reasoning module generates the next improvement hypothesis, serving as the gradient signal that directs optimization. It first extracts challenges $\mathcal{C}^{(i)}_{t}$ from structured feedback $f^{(i)}_{t}$ and local history $\mathcal{H}^{(i)}$, shifting from exploratory analysis (early iterations) to targeted diagnosis (addressing specific execution failures). For each identified challenge $c$, the module generates a concrete hypothesis $\eta_{c}$ by conditioning on the trace’s complete local state ($s^{*(i)},f^{(i)}_{t},\mathcal{H}^{(i)}$).
Candidates are scored across dimensions $\mathcal{D}$ (impact, alignment, novelty, feasibility, risk-reward) [^2] Weights: $(0.4,0.2,0.2,0.1,0.1)$ respectively.:
@@
where the success memory $\mathcal{M}$3.3) modulates scoring: hypotheses resembling past successes gain confidence, while similar failures receive penalties. Rather than greedily selecting the top-scored hypothesis, we sample from top-$k$ to maintain exploration diversity (in multi-trace settings, cross-trace mechanisms further guide selection, see §3.5). The selected hypothesis is finally implemented as solution $s^{(i)}_{t+1}$ via code generation and iterative refinement.
@@ 3.5 Multi-trace Optimization
Single-trace optimization risks local optima and cannot recover from poor early decisions. Gome employs $N$ parallel traces that synchronize via $\mathcal{M}$, enabling online knowledge sharing where trace $i$’s successful discoveries immediately inform traces $j\neq i$.
@@
To prevent redundant exploration (which renders memory sharing ineffective), we maximize initial search coverage by strictly enforcing diversity. Trace $n$’s starting hypothesis is conditioned on all prior proposals to ensure orthogonality:
$\eta_{0}^{(n)}=\texttt{GenerateHypothesis}\bigl(\mathcal{T},\{\eta_{0}^{(j)}\}_{j<n}\bigr),$ (6)
forcing the swarm to investigate distinct regions of the solution space from the outset.
@@
where $\{f_{\eta}\}$ denotes the associated feedback stored alongside each hypothesis. The selector has flexibility to select, modify, or generate based on the candidate pool and historical outcomes. Additional details of the probabilistic interaction kernel are provided in Appendix C.3.
@@ 4 Experiments
We evaluate Gome on MLE-Bench under a rigorous closed-world protocol to validate our claim that gradient-based optimization scales more effectively than search as reasoning capability improves. Additional experiments including protocol comparison and cost analysis are provided in Appendix A. Following prior MLE-Bench evaluations, we adopt any-medal rate as the primary metric. We additionally report improvement rate and validation-test correlation (IC) to characterize per-step optimization dynamics.
@@ 4.1 Experiment Setup
All experiments are conducted on MLE-Bench [chan2024mle-bench], a comprehensive benchmark comprising 75 Kaggle competitions for evaluating AI agents on machine learning engineering tasks, categorized by complexity into low, medium, and high. MLE-Bench-Lite consists of 22 low-complexity tasks for efficient evaluation. We report any-medal rate as the primary indicator.
Implementation.
We evaluate Gome using three frontier LLMs representing a spectrum of reasoning capabilities: DeepSeek-R1 [guo2025deepseek], o3 [openai_introducing_o3_2025], and GPT-5 [openai_introducing_gpt-5_2025], ordered by increasing reasoning strength. For the test environment, we use 12 vCPUs, 220GB RAM, and 1 NVIDIA V100 GPU with a 12-hour time budget. All results are averaged over three runs with different random seeds.
Baselines.
We benchmark against leading closed-world MLE agents on MLE-Bench, including MLAB [huang2023mlagentbench], OpenHands [wang2024openhands], AIDE [aide2025], AIRA [toledo2025ai], and ML-Master [liu2025ml]. For ML-Master, we re-evaluate under identical hardware and time budget to ensure fairness; for other baselines, we use results reported in their respective papers or the official leaderboard.
@@ 4.2 Main Results
Gome establishes a new state-of-the-art on MLE-Bench. As shown in Table 3, Gome outperforms all closed-world baselines, achieving a peak any-medal rate of 35.1% with GPT-5. Beyond aggregate performance, Gome excels in solution quality: it achieves a 96.0% valid submission rate and converts 16.4% of tasks to Gold medals. Under identical constraints (12h, V100), Gome matches ML-Master on weaker reasoners (DeepSeek-R1: 23.4% vs. 22.7%) but pulls ahead significantly with stronger models, widening the gap to +11.1 percentage points with GPT-5 and improving Gold rate. This widening gap is consistent with our hypothesis (validated across 10 models in Section 5). Furthermore, Gome matches the performance of AIRA despite operating under stricter resource constraints (half the time, weaker GPU), highlighting the efficiency of gradient-based optimization (full cross-protocol comparisons are provided in Table 8).
Accurate gradient signals drive rapid convergence. On MLE-Bench-Lite (Table 4), Gome achieves a 68.2% medal rate, matching the SOTA open-world method (Leeroo) despite lacking external retrieval. This confirms that for tractable tasks, internal diagnostics provide high-fidelity gradients sufficient for rapid convergence [nesterov2018lectures]. Conversely, the plateau on high-complexity tasks across all methods indicates a “noisy gradient” regime where reasoning limits are reached, suggesting that further gains require restoring signal accuracy through stronger models or external knowledge.
4.3 Ablation Study
To validate Gome’s design, we ablate its three core components: (1) w/o Structured Reasoning, replacing diagnostic analysis with simple error-correction prompts; (2) w/o Success Memory, removing the global repository of proven patterns; and (3) w/o Multi-trace Optimization, reducing to single-trajectory optimization without cross-trace knowledge sharing.
@@ 4.3 Ablation Study
As shown in Table 5, removing any component compromises the system’s overall robustness, particularly the primary Any-Medal rate. The Structured Reasoning ablation causes the most severe degradation, with improvement rate dropping from 41.1% to 22.6%, indicating that without diagnostic analysis, most optimization steps fail to produce valid improvements. Removing Success Memory shows a different failure mode: while per-iteration success remains decent, the lack of momentum from proven patterns leads to redundant exploration, reducing medal rate by 6.2%. Finally, Multi-trace Optimization is essential for diversity: its removal degrades final performance despite high local improvement rates, suggesting that cross-trace sharing helps the swarm escape the local optima that trap single-trace agents. All three components contribute meaningfully (Table 2).
@@ 5 Scaling Analysis
((a)) Scaling with Model Capability. As model capability increases from Efficiency to Frontier tiers, Gome’s advantage over search-based baselines widens significantly.
((b)) Convergence Dynamics. Gome (solid red) exhibits rapid early convergence, while MCTS (dashed blue) starts slower but catches up on weaker models. On Frontier-tier models, Gome maintains its advantage throughout.
Figure 3: Scaling analysis. (a) As reasoning capability improves, gradient-based optimization increasingly outperforms search baselines. (b) On weaker models, rapid early convergence is followed by a plateau due to noisy feedback; stronger reasoning enables sustained improvement throughout the 12h budget.
To validate our claim that gradient signal quality depends on reasoning capability, we evaluate Gome across three tiers: Efficiency (e.g., GPT-4o-mini), Advanced (e.g., DeepSeek-R1), and Frontier (e.g., o3, GPT-5) [^3] See Appendix A.1 for complete tier definitions and per-model results.. We compare against Gome-MCTS, a controlled variant replacing gradient updates with tree search within identical prompt and infrastructure, isolating the impact of the optimization strategy (implementation details in Appendix B.3).
Figure 3(a) reveals a distinct phase transition. On Efficiency-tier models, Gome lags behind: when gradient signals are noisy due to weak reasoning, exhaustive search outperforms directed updates. However, a crossover emerges as capability increases—search gains plateau while Gome accelerates, widening the gap to +5.8% on o3 and +7.1% on GPT-5. Figure 3(b) traces this divergence temporally. Gome exhibits rapid early convergence across all tiers, but sustainability differs: on weaker models (GPT-4o), unreliable gradients cause premature convergence, allowing MCTS to eventually catch up; on Frontier models, accurate gradient signals enable Gome to maintain its lead throughout, achieving both faster convergence and superior final performance. Complete per-model results and tier-wise gap visualization are provided in Appendix A.1.
@@ 6 Conclusion
We presented Gome, reframing MLE agents through gradient-based optimization. By treating LLM reasoning as gradient signals, success memory as momentum, and multi-trace execution as distributed optimization, Gome establishes a principled correspondence between classical optimization and agentic ML engineering.
Our experiments demonstrate state-of-the-art performance on MLE-Bench, while our scaling analysis reveals a more fundamental finding: tree search scales with inference compute (traversing more nodes), while Gome scales with model capability through increasingly accurate gradient signals. As foundation models advance in reasoning ability, this positions gradient-based optimization as an increasingly favorable paradigm.
This work opens a new design axis for MLE agents: rather than engineering more sophisticated search strategies, future systems may benefit from investing in gradient quality through richer feedback signals and stronger diagnostic reasoning. Future work may also explore hybrid strategies that combine tree search for early-stage exploration with gradient-based refinement as reasoning quality improves within a run. We release our codebase and GPT-5 execution traces to support reproducibility and future research.
@@
Gome’s effectiveness is sensitive to the base model’s reasoning capability. As shown in our scaling analysis, Gome underperforms tree-search methods on efficiency-tier models, where weaker reasoners may hallucinate diagnostic feedback. This also implies higher inference costs when using frontier models (Table 9). However, this dependency aligns with the trajectory of LLM development: as reasoning-oriented models advance rapidly, the capability threshold for effective gradient-based optimization will become increasingly accessible.
@@
We evaluated under a closed-world protocol on MLE-Bench, restricting agents to task-provided materials without external knowledge retrieval. While this ensures rigorous comparison of optimization strategies by controlling for confounding factors such as retrieval quality and knowledge source coverage, it limits generalizability: real-world MLE tasks often benefit from external knowledge, and Gome’s performance in open-world settings remains less thoroughly validated. Additionally, real Kaggle competitions allow participants to access shared notebooks and public baselines, so our closed-world setting may underestimate the performance achievable in practical scenarios.
@@
Our gradient-based framing establishes functional correspondence rather than mathematical equivalence. Unlike true gradients in continuous optimization, LLM-generated update signals lack formal properties such as convergence guarantees. This means that in rugged landscapes where solutions require fundamentally different paradigms rather than incremental edits, gradient-based guidance may get trapped in local optima. In practice, this is mitigated by the nature of MLE tasks: ML pipelines are typically repairable, making the smoothness assumption reasonable for most practical scenarios. Our multi-trace optimization and success memory further mitigate this risk by enabling shared exploration across trajectories.
@@ A.1 Extended Scaling Analysis
This section provides detailed per-model results supporting the scaling analysis in Section 5. We categorize models into three capability tiers following the stratification established by the Artificial Analysis Intelligence Index [^4] https://artificialanalysis.ai/, SWE-bench, and Aider leaderboard [^5] https://aider.chat/docs/leaderboards/ [artificialanalysis2026, jimenez2023swe, aider2026]: Efficiency models are optimized for throughput and cost; Advanced models represent balanced capability-efficiency trade-offs; Frontier models define current state-of-the-art reasoning.
@@ A.5 Task-type Breakdown
Table 10 presents detailed score comparison across 9 representative MLE-Bench tasks, grouped by data modality. We evaluate Gome with three frontier-tier backbone models: DeepSeek-R1 [guo2025deepseek], o3 [openai_introducing_o3_2025], and GPT-5 [openai_introducing_gpt-5_2025].
@@ B.1 Gome Hyperparameters
B.2 ML-Master Reproduction
We reproduce ML-Master [liu2025ml] as a representative state-of-the-art tree search baseline. We obtain the official implementation from the authors’ public repository [^6] https://github.com/sjtu-sai-agents/ML-Master and adapt it to our evaluation protocol.
@@ B.2 ML-Master Reproduction
Table 12: ML-Master reproduction validation on DeepSeek-R1. Original results are from [liu2025ml] on A100 GPUs; ours are averaged over 3 independent runs under closed-world protocol on V100 GPUs. All values are any-medal rates (%) on MLE-Bench with 12-hour budget.
@@ B.3 Gome-MCTS Implementation
Our modular framework naturally supports diverse exploration strategies beyond gradient-based optimization. To provide a fair comparison between optimization paradigms within the same framework, we implement Monte Carlo Tree Search (MCTS), a widely-adopted algorithm in existing MLE agents such as ML-Master [liu2025ml]. This implementation, called Gome-MCTS, serves two purposes: (1) it validates that our framework is a general-purpose platform capable of accommodating different search algorithms, and (2) it isolates the effect of the optimization strategy from other architectural differences when comparing against external baselines.